In 1993, the life expectancy of males in a certain country was 71.3 years. In 1999, it was 73.7 years. Let E represent the life expectancy in years t and let t represent the number of yers since 1993.
The linear function E(t) that fits the data is
E(t)=_t+_ (round to the nearest tenth)
Use the function to predict the life expectancy of males in 2004.
E(11)= (round to the nearest tenth)

Answer by stanbon(75887)

In 1993, the life expectancy of males in a certain country was 71.3 years.
In 1999, it was 73.7 years.
Let E represent the life expectancy in years t and let t represent the number of years since 1993.
--------
E(1993)= 71.3 is the y-intercept
-----------
Get the slope from (0,71.3) and (6,73.7)
m = (2.4/6) = 0.4
-----------
The linear function E(t) that fits the data is E(t)=0.4_t+71.3_ (round to the nearest tenth)
Use the function to predict the life expectancy of males in 2004.
E(11)= 0.04*11+71.3 = 75.7 (round to the nearest tenth)
============
Cheers,
Stan H.